(e) Suppose there is free entry and exit of firms in this industry.

The demand curve for a product is given by Q = 800 – 0.1P where P denotes price and Q denotes quantity of the product. The industry comprises large number of firms. Each firm's cost function is 

C(q) = 2000 + 500q+ 20q2


(e) Suppose there is free entry and exit of firms in this industry. Upon entry, the firms act as price-takers. Find the equilibrium price and aggregate output. How many firms will be there in this industry in the long run?
Answer #1

C= 2000 tsooq + zog? Aca 2000 + Soo + 20g d (Ac) = - 2000 + 20 = 0 - 2000 + 2092 = 0 q? - 100 q=10 qemin = 10 AC min = 2000 +